The Poseidon Swim Company produces swim trunks. The average selling price for one of their swim trunks is $78.78. The variable cost per unit is $25.52, Poseidon Swim has average fixed costs per year of $6,099. Determine the degree of operating leverage for the level of production and sales 342 swim trunks.

Round the answer to two decimal places.
Answer #1
Sales(78.78*342) 26942.76
Less:Variable cost(25.52*342) 8727.84
Contribution margin $18214.92
Less:Fixed Costs 6099
Operating income $12115.92

Degree of operating leverage=Contribution margin/Operating income

=$18214.92/$12115.92

which is equal to

=1.50(Approx).
Answer #2

The degree of operating leverage (DOL) is a measure of how sensitive a company's operating income (or profit) is to changes in its sales revenue. It can be calculated using the following formula:

DOL = (Contribution Margin / Operating Income)

Where: Contribution Margin = (Selling Price per Unit - Variable Cost per Unit) Operating Income = (Total Sales - Total Variable Costs - Total Fixed Costs)

Let's calculate the degree of operating leverage for Poseidon Swim Company:

Given data: Selling Price per Unit = $78.78 Variable Cost per Unit = $25.52 Total Fixed Costs = $6,099 Number of swim trunks produced and sold = 342

Contribution Margin = (78.78 - 25.52) = $53.26

Total Sales Revenue = (Number of units sold × Selling Price per Unit) Total Sales Revenue = (342 × 78.78) = $26,912.76

Total Variable Costs = (Number of units sold × Variable Cost per Unit) Total Variable Costs = (342 × 25.52) = $8,720.64

Operating Income = (Total Sales Revenue - Total Variable Costs - Total Fixed Costs) Operating Income = (26,912.76 - 8,720.64 - 6,099) = $12,093.12

Now, calculate the degree of operating leverage (DOL): DOL = (53.26 / 12,093.12)

DOL ≈ 0.0044

Rounded to two decimal places, the degree of operating leverage for the production and sales of 342 swim trunks is approximately 0.00.

answered by: Hydra Master